TOWARDS A MORI THEORY ON COMPACT KÄHLER THREEFOLDS, II

THOMAS PETERNELL

Abstract. This paper, a continuation of “Towards a Mori theory on compact Kähler manifolds, 1” (written with F. Campana), introduces a non-algebraic analogue to Mori theory in dimension 3.